### v1

# PAIR-Former: Budgeted Relational Multi-Instance Learning for Functional miRNA Target Prediction

Abstract
Functional miRNA–mRNA targeting is a large-bag prediction problem: each transcript yields a heavy-tailed pool of candidate target sites (CTSs), yet only a pair-level label is observed. We formalize this regime as Budgeted Relational Multi-Instance Learning (BR-MIL), where at most $K$ instances per bag may receive expensive encoding and relational processing under a hard compute budget. We propose PAIR-Former (Pool-Aware Instance-Relational Transformer), a BR-MIL pipeline that performs a cheap full-pool scan, selects up to $K$ diverse CTSs on CPU, and applies a permutation-invariant Set Transformer aggregator on the selected tokens. On miRAW, PAIR-Former outperforms strong pooling baselines at a practical operating budget ($K^{\star}{=}64$) while providing a controllable accuracy–compute trade-off as $K$ varies. We further provide theory linking budgeted selection to (i) approximation error decreasing with $K$ and (ii) generalization terms governed by $K$ in the expensive relational component.
miRNA-mRNA target prediction, multi-instance learning, budgeted inference, instance selection, permutation-invariant aggregation
Code availability.
An anonymized implementation is included in the supplementary material; we will release the full codebase upon acceptance.
Figure 1: Overview of PAIR-Former under BR-MIL. Top: Inference pipeline. Candidate CTSs are generated on the $3^{\prime}\mathrm{UTR}$ using a TargetNet-compatible ESA scan/filter ($s_{i}^{\mathrm{esa}}\geq 6$), producing a variable-size heavy-tailed bag with $n$ candidates. A cheap student encoder $\tilde{e}_{\tilde{\theta}}$ scans all $n$ instances to obtain cheap logits $\tilde{z}_{i}$ and embeddings $\tilde{h}_{i}$. Under a strict per-bag budget $K=\min(\texttt{kmax},n)$, STSelector selects a subset $S$ ($|S|\leq K$) by combining Top-$K_{1}$ exploitation (by $\tilde{z}_{i}$) and diversity over transcript position $p_{i}$ and embedding redundancy. Only selected sites are processed by the expensive encoder $e_{\theta}$ to form CTS tokens (e.g., concatenating $h_{i}$, $z_{i}$, $s_{i}^{\mathrm{esa}}$, and $p_{i}$), padded/masked to kmax, and aggregated by a permutation-invariant Set Transformer (SAB stack + PMA) into a pair logit $z_{\text{pair}}$ and prediction $\hat{Y}$. Bottom: Three-stage training: (1) train the expensive CTS encoder $e_{\theta}$ with window-level supervision; (2) distill the cheap encoder $\tilde{e}_{\tilde{\theta}}$ from a teacher; (3) train the pair-level aggregator $f_{\phi}$ using the budgeted forward, with an encoder-freeze warmup followed by joint fine-tuning.
1 Introduction
MicroRNAs (miRNAs) are short non-coding RNAs that regulate gene expression by binding to partially complementary sites on messenger RNAs (mRNAs), most commonly within 3 untranslated regions ($3^{\prime}\mathrm{UTR}s$). A central computational task is functional miRNA–mRNA interaction prediction: given a miRNA and a candidate transcript, predict whether the transcript is functionally repressed (binary classification at the transcript/pair level). Compared with detecting sequence complementarity alone, functional prediction is harder because the observed label reflects downstream regulatory outcomes and thus aggregates multiple latent factors rather than providing instance-level supervision at individual binding sites (Bartel, 2009; Min et al., 2022).
Most modern pipelines follow a common decomposition: (i) generate candidate target sites (CTSs) on the $3^{\prime}\mathrm{UTR}$ via heuristic rules or relaxed alignment/seed criteria, (ii) score each miRNA-CTS window independently using a window-level predictor, and (iii) aggregate site scores into a transcript-level prediction, frequently via max pooling (Min et al., 2022; Pla and others, 2018; Li and others, 2024; Yang et al., 2024). While window scorers have improved substantially, the final aggregation step typically encodes a strongest-site assumption: the bag label is driven by the single most confident site, and multiple CTSs are otherwise treated as conditionally independent given the miRNA.
This inductive bias can be limiting when a transcript contains many candidates with redundancy, cooperation, or competition among sites. Crucially, the number of candidate sites varies widely across transcripts and can be heavy-tailed, making naive relational aggregation over all sites computationally infeasible in practice.
We therefore cast functional miRNA–mRNA prediction as multi-instance learning (MIL): each miRNA–mRNA pair forms a bag, its CTS windows are instances, and only the bag label is observed (Carbonneau et al., 2018). However, naively upgrading the aggregator from max pooling to a relational model (e.g., self-attention over all instances) is computationally prohibitive in this domain: the candidate pool size $n$ is large and varies substantially across pairs (often heavy-tailed), and full self-attention would require $\mathcal{O}(n^{2})$ time/memory. This motivates a compute-aware formulation that explicitly separates (i) cheap per-instance scanning from (ii) expensive relational reasoning.
Budgeted Relational MIL.
We formalize this setting as Budgeted Relational Multi-Instance Learning (BR-MIL): each example is a labeled bag with a large candidate pool, but the model may apply expensive encoding and relational aggregation to at most $K$ instances per bag. A cheap encoder scans all candidates, a selector chooses $|S|\leq K$, and a permutation-invariant aggregator predicts from the selected tokens, reducing the dominant relational cost from $\mathcal{O}(n^{2})$ to $\mathcal{O}(K^{2})$.
Our method: PAIR-Former.
We propose PAIR-Former (Pool-Aware Instance-Relational Transformer), a practical BR-MIL instantiation for functional miRNA targeting. To preserve comparability with established protocols, we keep candidate generation and alignment-aware encoding TargetNet-compatible (Min et al., 2022; Li and others, 2024) and focus innovation on budgeted relational reasoning: (i) we train a strong CTS-level “teacher” encoder and distill a lightweight “student” that can cheaply scan all candidates; (ii) we introduce a CPU-only budgeted selector (STSelector) that combines Top-$K$ exploitation with diversity over transcript position and cheap embedding redundancy; (iii) we apply an expensive encoder only on the selected $K$ CTSs and aggregate them using a permutation-invariant Set Transformer (self-attention blocks) to capture redundancy/cooperation/competition patterns without relying on input order.
Theory (high level).
We show that budgeted execution can be viewed as a controlled approximation of a full-information relational predictor, and that the dominant generalization term of the expensive relational component scales with $K$ rather than the raw bag size $n$ (statements in Sec. 5).
Contributions.
  • Problem formulation (BR-MIL). We formalize functional miRNA–mRNA prediction as Budgeted Relational MIL, capturing large/heavy-tailed candidate pools and a strict per-bag budget $K$ for expensive relational reasoning, while enforcing permutation-invariant prediction.
  • Method (PAIR-Former). We develop a practical BR-MIL pipeline with (i) TargetNet-compatible CTS generation/encoding, (ii) teacher–student distillation for cheap full-pool scanning, (iii) a diverse budgeted selector (STSelector), and (iv) a permutation-invariant Set Transformer aggregator operating on $K$ selected tokens.
  • Theory–experiment evidence chain. We provide two theorems linking the budget $K$ to (i) approximation error under truncation and (ii) generalization of the expensive relational component, and empirically validate them via controlled $K$-sweeps (Fig. 2), runtime scaling (Fig. 3), and robustness to candidate pool size $n$ (Fig. 5).
2 Related Work
Functional miRNA target prediction.
Classical functional miRNA target predictors rely on hand-crafted signals that emphasize seed complementarity, conservation, and biophysical/context features (e.g., accessibility and thermodynamic stability), as exemplified by TargetScan, miRanda, PITA, mirSVR, and miRDB (Lewis et al., 2005; Enright et al., 2003; Kertesz et al., 2007; Betel et al., 2010; Wong and Wang, 2015). These approaches are lightweight and often interpretable, but can be brittle when interaction patterns deviate from canonical assumptions (Bartel, 2009; Grimson et al., 2007). Modern deep models reduce manual feature engineering by learning sequence/alignment-driven representations (Wen and others, 2016; Pla and others, 2018). A widely used site-first pipeline enumerates candidate target sites (CTSs), scores each miRNA–CTS window independently, and aggregates site scores into a transcript-level prediction, often via max pooling (Min et al., 2022; Pla and others, 2018; Li and others, 2024; Yang et al., 2024). Despite substantial progress on window-level encoders (e.g., TargetNet-style alignment-aware encodings and stronger backbones), transcript-level inference in many pipelines still operationalizes a strongest-site assumption and largely treats sites as conditionally independent given the miRNA (Min et al., 2022; Sætrom et al., 2007).
Multi-instance learning and set-based aggregation.
The weakly supervised nature of functional targeting naturally aligns with multi-instance learning (MIL), where only bag-level labels are observed and instance labels are latent (Dietterich et al., 1997; Carbonneau et al., 2018). Permutation-invariant set modeling provides a principled interface for MIL: DeepSets-style pooling offers simple inductive bias via sum/mean/max (Zaheer et al., 2017), attention-based MIL learns data-dependent instance weights under bag supervision (Ilse et al., 2018), and Set Transformer extends this by modeling higher-order interactions among instances through self-attention while preserving permutation invariance (Lee et al., 2019). Biologically, multiple sites within a $3^{\prime}\mathrm{UTR}$ can exhibit redundancy, cooperation, or competition, and transcript-level outcomes can be modulated by alternative targets/decoy-like binding (Bartel, 2009; Sætrom et al., 2007; Salmena et al., 2011), motivating relational aggregation beyond max pooling.
Positioning: budgeted relational MIL for heavy-tailed candidate pools.
A key practical challenge in miRNA targeting is that the CTS pool size is large and heavy-tailed across transcripts, making naive relational aggregation over all instances computationally infeasible. Unlike standard attention-based MIL or Set Transformer formulations that implicitly assume all instances can be encoded and attended to, BR-MIL enforces a hard per-bag budget $K$ on expensive encoding and relational computation. PAIR-Former instantiates this regime by combining (i) a cheap full-pool scan, (ii) budgeted and diversity-aware subset selection, and (iii) permutation-invariant Set Transformer aggregation on the selected tokens, while keeping candidate generation and alignment-aware CTS encoding TargetNet-compatible for protocol comparability (Min et al., 2022; Li and others, 2024). This design targets the gap between strong site scorers and realistic compute constraints in large-bag functional targeting.
3 Problem Setup
Large-bag functional targeting.
Each miRNA–mRNA pair defines a bag of candidate target sites (CTSs): $B=\{(x_{i},u_{i})\}_{i=1}^{n}$, where $x_{i}$ is the CTS content (e.g., alignment-aware window tensor) and $u_{i}$ are structural attributes (e.g., normalized transcript position and ESA score). The pool size $n$ is large and heavy-tailed across transcripts, and only a pair-level label $Y\in\{0,1\}$ is observed.
BR-MIL (one-sentence).
BR-MIL is multi-instance learning with heavy-tailed candidate pools, where expensive encoding and relational aggregation may be applied to at most $K$ instances per bag under a hard compute budget.
Budgeted predictor, permutation invariance, and cost.
A cheap encoder $\tilde{e}_{\tilde{\theta}}$ scans all $n$ candidates to produce cheap signals (e.g., $\tilde{h}_{i},\tilde{z}_{i}$), a selector $\pi_{\psi}$ returns a subset $S(B)\subseteq[n]$ with $|S(B)|\leq K$, and only selected instances are expensively encoded and aggregated by a permutation-invariant set function:
$\begin{split}\hat{Y}(B)&=f_{\phi}\Big(\big\{\tau(e_{\theta}(x_{i},u_{i}),u_{i})\big\}_{i\in S(B)}\Big),\\ S(B)&=\pi_{\psi}\big(\{\tilde{e}_{\tilde{\theta}}(x_{i},u_{i})\}_{i=1}^{n}\big).\end{split}$ (1)
Permutation invariance requires $\hat{Y}(B)=\hat{Y}(\pi(B))$ for any permutation $\pi$ of the instances, which holds when selection depends only on the multiset of cheap signals (with fixed tie-breaking) and $f_{\phi}$ is permutation-invariant (e.g., Set Transformer / DeepSets). Per bag, the pool-dependent work is $\mathcal{O}(n)$ (cheap scan/selection) while the expensive relational component is governed by the budget, typically $\mathcal{O}(K)$ for expensive encoding and $\mathcal{O}(K^{2})$ for attention-based aggregation. Full notation, the formal BR-MIL definition/objective, and invariance conditions are deferred to Appendix A.
4 Method
4.1 Overview: BR-MIL with budgeted selection and relational aggregation
Given a miRNA–mRNA pair $(\mu,\nu)$ with transcript-level label $Y\in\{0,1\}$, we extract a variable-size pool of candidate target sites (CTSs) from the 3’UTR, yielding a bag $B=\{(x_{i},u_{i})\}_{i=1}^{n}$ where $x_{i}$ is an alignment-aware CTS content tensor and $u_{i}$ contains structural attributes (position and alignment score). Our BR-MIL instantiation enforces a strict per-bag budget $K\leq 64$ on expensive instance encoding and relational aggregation via a three-component pipeline: (i) a cheap instance encoder scans all $n$ candidates to produce cheap scores/embeddings; (ii) a CPU-only STSelector selects a diverse subset $S\subseteq[n]$ with $|S|=K$; (iii) an expensive instance encoder is evaluated only on $S$, and the resulting CTS tokens are aggregated by a Set Transformer (SAB) into a pair-level logit $z_{\text{pair}}$ and prediction $\hat{Y}=\sigma(z_{\text{pair}})$. An end-to-end overview of candidate generation, budgeted selection, permutation-invariant aggregation, and the three-stage training recipe is shown in Fig. 1.
4.2 CTS generation and ESA encoding (TargetNet-compatible)
CTS extraction via ESA filtering.
We scan the 3’UTR $\nu$ using a 40-nt sliding window with stride 1. For each window, we compute an extended seed alignment (ESA) score $s_{i}^{\mathrm{esa}}$ between the miRNA extended seed region and the window, and retain candidates with $s_{i}^{\mathrm{esa}}\geq 6$ (a relaxed threshold that admits both canonical and non-canonical patterns while filtering unstable candidates).
Alignment-aware tensor encoding.
For each retained candidate, we construct an ESA-based input tensor $X_{i}\in\mathbb{R}^{10\times 50}$ following the TargetNet/miTDS protocol (kept unchanged for comparability). We use $x_{i}:=X_{i}$ as the instance content.
Structural attributes.
We use two structural attributes: the length-normalized CTS center position $p_{i}=\operatorname{center}(c_{i})/|\nu|\in[0,1].$ and the ESA score $s_{i}^{\mathrm{esa}}$. We denote $u_{i}:=(p_{i},s_{i}^{\mathrm{esa}})$.
4.3 Instance encoders: expensive teacher and cheap student
We instantiate two CTS encoders:
$(h_{i},z_{i})$ $=e_{\theta}(x_{i},u_{i}),\qquad h_{i}\in\mathbb{R}^{384},~z_{i}\in\mathbb{R},$ (2)
$(\tilde{h}_{i},\tilde{z}_{i})$ $=\tilde{e}_{\tilde{\theta}}(x_{i},u_{i}),\qquad\tilde{h}_{i}\in\mathbb{R}^{64},~\tilde{z}_{i}\in\mathbb{R}.$ (3)
The expensive encoder $e_{\theta}$ is TargetNet-style and augmented with channel attention while preserving input/output compatibility; the cheap encoder $\tilde{e}_{\tilde{\theta}}$ is trained by distillation from $e_{\theta}$ (Sec. 4.7).
4.4 Budgeted subset selection with STSelector
For each bag with $n$ candidates, we set $K=\min(\texttt{kmax},n)$ (default $\texttt{kmax}=64$) and split the budget into $K_{1}=\lfloor\rho K\rfloor$ and $K_{2}=K-K_{1}$ with $\rho=0.5$.
Design principle.
Under heavy-tailed candidate pools, pure Top-$K$ selection tends to concentrate on near-duplicate sites. We therefore design STSelector to combine: (i) exploitation of high-confidence candidates under the cheap encoder, and (ii) coverage/diversity over transcript position and embedding space, while remaining CPU-only for low-latency inference.
This design is aligned with the approximation view in Sec. 5: by spreading the budget over high-scoring yet non-redundant candidates, the selector aims to cover a larger fraction of the total influence mass.
High-level behavior.
STSelector first selects a Top-$K_{1}$ subset according to cheap logits $\tilde{z}_{i}$, and then fills the remaining $K_{2}$ slots by encouraging coverage across transcript positions and reducing redundancy in the cheap embedding space. The output is a diverse subset $S\subseteq[n]$ with $|S|=K$ that preserves high-scoring candidates while avoiding over-concentration.
Complexity.
Selection runs in near-linear time in $n$ (dominated by lightweight per-bin top-$m$ operations), and does not require GPU execution.
Implementation details.
Full algorithmic steps (binning, SimHash-based deduplication, and balanced quota allocation) are provided in Appendix D.
4.5 Tokenization and Set Transformer (SAB) aggregator
Per-CTS token.
After selection, we evaluate the expensive encoder only on $i\in S$ and define a hard token for each selected CTS by concatenation:
$t_{i}~=~\big[h_{i}~\|~z_{i}~\|~s_{i}^{\mathrm{esa}}~\|~p_{i}\big]~\in~\mathbb{R}^{387}.$ (4)
We pad to length kmax with a binary mask so that padding positions do not participate in attention/aggregation.
Relational aggregator (Set Transformer, SAB).
Let $T=\{t_{i}\}_{i\in S}$ be the masked token set. We use a Set Transformer stack with self-attention blocks (SAB), where the canonical definition is $\mathrm{SAB}(X)=\mathrm{MAB}(X,X)$, i.e., full self-attention over the set. The aggregator outputs a pair-level logit $z_{\text{pair}}\in\mathbb{R}$ and prediction $\hat{Y}=\sigma(z_{\text{pair}})$. Self-attention implicitly captures CTS–CTS interaction effects (competition/redundancy/cooperation) without explicitly constructing relation features. By construction, the aggregator only observes the selected subset and is permutation invariant with respect to its inputs, satisfying the BR-MIL invariance requirement in Sec. 3.
4.6 Three-stage training procedure
We train the pipeline in three stages, aligning supervision level and compute budget:
Stage 1: train expensive CTS model ($e_{\theta}$).
We train the instance encoder $e_{\theta}$ on miRNA–CTS pairs with window-level binary supervision (TargetNet-compatible protocol).
Stage 2: distill cheap CTS model ($\tilde{e}_{\tilde{\theta}}$).
We train $\tilde{e}_{\tilde{\theta}}$ on the same miRNA–CTS data by distilling from the expensive teacher, using a composite distillation objective combining supervised loss, logit distillation, feature matching, and (optional) relational distillation (Sec. 4.7).
Stage 3: train Set Transformer aggregator ($f_{\phi}$) and joint fine-tune.
We train the aggregator on miRNA–mRNA pairs using the strict budgeted forward: CTS scan $\rightarrow$ cheap scan over all $n$ $\rightarrow$ STSelector selects $K$ $\rightarrow$ expensive encoding on $K$ $\rightarrow$ tokenization $\rightarrow$ Set Transformer $\rightarrow$ pair loss. We use a two-phase strategy: (i) warmup: freeze the instance encoder and train only the aggregator; (ii) joint fine-tune: unfreeze the instance encoder and fine-tune both modules for a small number of epochs to reach fast convergence.
4.7 Losses
We use a unified binary classification objective for both CTS-level training and pair-level aggregation, and train the cheap encoder by teacher–student distillation.
Binary classification.
Given a logit $z$ and label $y\in\{0,1\}$, we use a numerically stable binary loss (BCE with logits), optionally combined with label smoothing and focal reweighting to handle class imbalance and hard examples.
Distillation for the cheap encoder.
The cheap encoder $\tilde{e}_{\tilde{\theta}}$ is trained to approximate the expensive teacher $e_{\theta}$ using a composite objective that combines: (i) supervised loss on student logits, (ii) temperature-scaled logit distillation, (iii) feature-level matching between teacher and student embeddings, and (iv) optional relational distillation via similarity matrices.
$\begin{split}L_{\mathrm{distill}}=(1-\alpha)\,L_{\mathrm{sup}}&+\alpha\,L_{\mathrm{KD}}\\ &+\beta_{\mathrm{feat}}\,L_{\mathrm{feat}}+\beta_{\mathrm{rel}}\,L_{\mathrm{rel}}\,.\end{split}$ (5)
Pair-level loss.
The Set Transformer aggregator outputs a single pair logit $z_{\mathrm{pair}}$ per miRNA–mRNA example, which is trained using the same binary loss family.
Hyperparameters and exact formulations.
Full loss definitions, weighting coefficients, and schedules are provided in Appendix E.
4.8 Inference
At test time, inference for one miRNA–mRNA pair follows: (1) CTS scan (40-nt, stride 1) + ESA filtering ($s^{\mathrm{esa}}\geq 6$) to obtain $n$ candidates; (2) cheap scan over all $n$ to get $(\tilde{z}_{i},\tilde{h}_{i})$; (3) STSelector selects $S$ with $|S|=K=\min(\texttt{kmax},n)$; (4) expensive encoding only on $i\in S$ to get $(h_{i},z_{i})$; (5) tokenize via Eq. (4) with padding/mask to kmax; (6) Set Transformer (SAB) $\rightarrow z_{\text{pair}}\rightarrow\hat{Y}=\sigma(z_{\text{pair}})$. Pseudocode for the training and inference pipelines is provided in Appendix C.
5 Theory: Budgeted Relational MIL
We summarize two results that formalize BR-MIL as: (i) a controlled approximation of a full-information relational predictor under a budget constraint, and (ii) a hypothesis class whose dominant generalization term for the expensive relational component depends on $K$ rather than the raw bag size $n$.
5.1 Approximation under budgeted expensive encoding
Consider a reference predictor that applies the same relational aggregator to all expensive tokens in a bag, and its budgeted counterpart that runs the aggregator only on a selected subset $S$ with $|S|\leq K$ (using a masking convention).
Theorem 5.1 (Approximation error vs. budget $K$).
Under mild stability and boundedness assumptions, the expected prediction gap between the full-information predictor and its budgeted counterpart is upper bounded by a constant times the uncovered influence mass of unselected instances. In particular, the approximation error decreases as $K$ increases and as the selector improves coverage of influential instances.
This result formalizes budgeted execution as a controlled truncation of relational computation: increasing $K$ reduces the mass of discarded information and tightens the approximation.
5.2 Generalization controlled by $K$
We next characterize the statistical behavior of the expensive relational component in BR-MIL.
Theorem 5.2 (Generalization governed by budget $K$).
Let $f$ be a permutation-invariant relational aggregator that receives at most $K$ selected tokens per bag. Conditioned on a fixed selector, standard uniform-convergence arguments yield a dominant generalization term scaling as $\tilde{\mathcal{O}}(\sqrt{K/M})$ in the number of bags $M$, independent of the raw candidate pool size $n$. Any additional dependence on $n$ can only enter through the selector family.
Together, these results motivate the BR-MIL structure: a cheap scan over all candidates, budgeted selection of at most $K$ instances, and permutation-invariant relational aggregation only on the selected subset.
Proofs and assumptions.
Full assumptions and proofs of Theorems 5.1 and 5.2 are provided in Appendix B.
6 Experiments
6.1 Datasets and evaluation protocol
We evaluate functional miRNA–mRNA interaction prediction using the public resources released with TargetNet/miTDS (Min et al., 2022; Li and others, 2024). In particular, the release provides (i) a miRAW-derived miRNA–CTS dataset for window-level supervision, and (ii) ten predefined, class-balanced miRAWtest subsets for pair-level evaluation, each containing 548 positive and 548 negative miRNA–mRNA pairs (Pla and others, 2018; Min et al., 2022).
Stage 1–2 (CTS-level) data.
We train the expensive CTS encoder (Stage 1) and distill the cheap CTS encoder (Stage 2) using the released miRNA–CTS train/validation split provided by TargetNet (CTS-level supervision), following their preprocessing pipeline (Min et al., 2022; Li and others, 2024). These stages use only CTS-level labels and do not use any pair-level labels.
Stage 3 (pair-level) data: released miRAWtest half-split.
The full large-scale miRAW pair corpus used in the original TargetNet protocol is not fully available as a public training split. Therefore, to obtain a reproducible pair-level training signal for the aggregator, we construct a released miRAWtest half-split using only the ten released miRAWtest subsets. Concretely, we partition the ten subsets into two disjoint halves: we use subsets $\{1,2,3,4,5\}$ (5,480 pairs) as a development pool and split it into train/validation with a 0.9/0.1 ratio, while reserving subsets $\{0,6,7,8,9\}$ (5,480 pairs) as a held-out test partition. This split is fixed, uses no selective sampling, and preserves class balance by construction.
Protocol caveat (comparability).
Because our Stage-3 pair-level training is performed on a subset of the released miRAWtest pairs, our absolute numbers are not directly comparable to results reported under the full miRAW pair protocol in prior work. We therefore (i) reproduce key baselines under the identical half-split and identical evaluation code, and (ii) interpret quoted results from prior work as context only.
Metrics.
We report PR-AUC (AUPR) as the primary metric due to its robustness for ranked prediction and its prevalence in recent functional miRNA target prediction evaluations (Min et al., 2022; Yang et al., 2024). As a complementary thresholded metric, we report $\mathrm{F1}@0.5$ using a fixed decision threshold of 0.5. Optionally, we also report $\mathrm{F1}_{\text{best}}$ by selecting the threshold that maximizes F1 on the validation set and applying it to the test set (reported where noted).
6.2 Baselines and model variants
Reproduced deep baselines on our half-split.
We re-run TargetNet (Min et al., 2022) under our half-split and evaluate it using the same metric computation code as our methods.
Quoted baselines (context only).
For miTDS (Li and others, 2024) and representative classic baselines (e.g., PITA), we cite numbers reported in prior work. These results are typically reported under their original evaluation setting (often aggregating over all ten released subsets), and are therefore not directly comparable to our held-out half-split. We include them only for context and focus our claims on reproduced baselines under the identical split and evaluation code.
Our main method and matched internal baseline.
Our primary comparison is between: (i) a strong TargetNet-style encoder + max pooling baseline (k-independent at inference), and (ii) BR-MIL with a fixed operating budget $K^{\star}=\mathbf{64}$ throughout the main paper. BR-MIL uses cheap signals to select a size-$K$ subset of candidates and applies an expensive encoder and a permutation-invariant relational aggregator only on the selected subset (see Sec. 4).
6.3 Implementation details
Default operating budget.
Unless otherwise specified, we use $K^{\star}=\mathbf{64}$ as the default operating point for BR-MIL and all main comparisons (Table 1, Fig. 23, and Fig. 5). For selector ablations, we additionally report $K\in\{32,64\}$ in Table 2.
Training/evaluation repeats.
Unless otherwise noted, we run $R$ independent trials with seeds $\{2020,2025,2026\}$ (thus R=3R{=}3), select checkpoints by validation PR-AUC, and report test metrics from the selected checkpoint. Test subsets are never used for early stopping, threshold selection, or any form of caching across splits; all caches (cheap scan, selection indices, and packed tensors) are computed and stored separately for train/validation vs. test partitions.
Hardware.
All experiments are conducted on a single NVIDIA RTX 4090 GPU with 24 GB memory. We use PyTorch 2.4.1 (+cu121) with CUDA 12.1 and cuDNN 9.1.0.
6.4 Main results
Table 1: Overall performance on the released miRAWtest half-split (held-out test subsets). We report mean performance for reproduced baselines and our methods under the same split and the same evaluation code. Quoted baselines are taken from prior work under their original evaluation settings and are shown for context only (not directly comparable).
Method PR-AUC F1@0.5 Acc Prec Rec Spec NPV
Reproduced baselines (miRAW half-split)
TargetNet (reproduced) (Min et al., 2022) 0.7632 0.7714 0.7226 0.6560 0.9361 0.5091 0.8885
Ours (miRAW half-split)
Max pooling (ours, expensive encoder + max pooling) 0.8264 0.8125 0.8049 0.7822 0.8453 0.7646 0.8317
PAIR-Former (ours) 0.9961 0.9740 0.9744 0.9886 0.9598 0.9890 0.9612
Quoted baselines (full miRAW protocol; not strictly comparable)
miTDS (quoted) (Li and others, 2024) 0.8063 0.7700 0.6962 0.9578 0.5821 0.9326
PITA (quoted) (Kertesz et al., 2007) 0.2162 0.5053 0.5196 0.1365 0.8741 0.5030
miRDB (quoted) (Wong and Wang, 2015) 0.2110 0.5373 0.7135 0.1239 0.9507 0.5205
miRanda (quoted) (Enright et al., 2003) 0.3568 0.5001 0.4997 0.2775 0.7226 0.5001
TargetScan (quoted) (Lewis et al., 2005) 0.4712 0.5577 0.5852 0.3945 0.7208 0.5436
deepTarget (quoted) (Wen and others, 2016) 0.4904 0.6521 0.8332 0.3477 0.9354 0.6064
miRAW (quoted) (Pla and others, 2018) 0.7289 0.7055 0.6749 0.7923 0.6186 0.7493
Main result.
Under the released miRAWtest half-split and a fixed operating budget $K^{\star}{=}64$, PAIR-Former achieves strong performance (PR-AUC $0.9961$, F1@0.5 $0.9740$), outperforming TargetNet and a matched max-pooling baseline. We note that miRAWtest subsets are small and class-balanced, so absolute metrics may saturate; we focus on improvements under the identical split and evaluation code.
6.5 Performance vs. budget $K$
[…]
Link to Theorem 5.1.
Figure 2 isolates truncation error by evaluating a fixed $K_{\max}$ model under smaller budgets via masking, and shows monotone gains as $K$ increases. Unless stated otherwise, we use the budget-matched setting with $K^{\star}{=}64$.
6.6 Runtime and memory under a fixed token budget $K$
We benchmark online inference cost as a function of the token budget $K$ and compare BR-MIL_online with TargetNet_like_online (budget-independent) and a heavier Naive_online variant. Figure 3 shows that at the operating point K=64K{=}64, BR-MIL achieves TargetNet-like end-to-end latency/throughput while enabling relational aggregation, whereas Naive is consistently slower. Peak VRAM grows with $K$ for budgeted relational pipelines but remains modest at $K\leq 64$. Figure 4 further indicates that overall wall-time is dominated by a shared CPU gather step; BR-MIL-specific selection and aggregation add only a small overhead at K=64K{=}64. Full profiling setup, absolute numbers, and component definitions are provided in Appendix H.
Figure 3: Online inference cost vs. budget $K$. End-to-end latency, throughput, and peak VRAM for BR-MIL_online and Naive_online across $K$, with TargetNet_like_online as a budget-independent reference. Bottom: representative profiled components.
Figure 4: Stage-wise latency breakdown at K=64K{=}64 (log scale). Across pipelines, the shared CPU gather stage dominates wall-time; BR-MIL selection/aggregation contribute a small overhead. See Appendix H for stage definitions and profiling details.
6.7 Selector ablation
[…]
Variant K=32K{=}32 K=64K{=}64
[…]
Ablation note.
The best selector variant depends on the budget. At K=32K{=}32, S2 improves over S0 by reducing redundancy via embedding deduplication. At K=64K{=}64, S1 is slightly better, suggesting that extra pruning can be overly aggressive when the budget is larger. We use S2 as the default due to its stronger small-budget performance and redundancy control in heavy-tailed pools.
6.8 Robustness to candidate pool size $n$
Figure 5: Robustness to visible candidate pool size $n$ at fixed budget $K^{\star}{=}64$ (evidence for Theorem 5.2). We cap the selector’s visible pool to the top-$n$ candidates ranked by the cheap logit $\tilde{z}_{i}$, with $n\in\{64,128,256,512,1024,2048\}$, while keeping the expensive-token budget and Set Transformer input length fixed at $K^{\star}$. We report test (a) PR-AUC and (b) $\mathrm{F1}@0.5$. Performance drops when $n\approx K^{\star}$ but saturates once $n$ is a few multiples of $K^{\star}$ (e.g., $n\geq 256$), suggesting that varying $n$ mainly affects selection quality rather than the capacity of the expensive relational component.
Observation.
Performance saturates once $n$ is a few multiples of $K^{\star}$ (e.g., $n\geq 256$), consistent with Theorem 5.2.
7 Discussion
Takeaways.
Functional miRNA–mRNA targeting exhibits heavy-tailed candidate pools, making full relational aggregation over all CTSs impractical. BR-MIL explicitly separates cheap full-pool scanning from budgeted expensive relational reasoning, so the dominant relational cost is governed by $K$ rather than raw pool size $n$. Within this framework, PAIR-Former replaces strongest-site pooling with (i) budgeted, diversity-aware selection and (ii) permutation-invariant Set Transformer aggregation, capturing cross-site redundancy/competition/cooperation while preserving TargetNet-compatible candidate generation for protocol comparability.
Limitations and future work.
Our supervision is primarily pair-level; thus, learned instance importance/attention should not be interpreted mechanistically, and performance is bounded by label noise and curation artifacts. BR-MIL also inherits dependence on candidate generation: unretrieved functional sites are irrecoverable downstream. Hard selection under small budgets can miss weak-but-combinatorial evidence; more principled budgeted selection (e.g., stochastic/differentiable subset selection with exchangeable noise) is an important direction. Finally, Set Transformer attention scales as $\mathcal{O}(K^{2})$; sparse/linear attention or hierarchical pooling could extend BR-MIL to larger budgets and broader functional datasets beyond miRAW.
8 Conclusion
We formulated functional miRNA–mRNA targeting as BR-MIL, where heavy-tailed candidate pools require a strict per-bag budget $K$ for expensive encoding and relational aggregation. We proposed PAIR-Former, which performs a cheap full-pool scan, selects a diverse size-$K$ subset, and applies a permutation-invariant Set Transformer on the selected CTS tokens. On miRAW, PAIR-Former substantially improves over pooling baselines at a practical operating budget ($K^{\star}{=}64$) while yielding a controlled accuracy–compute trade-off as $K$ varies. Our theory and controlled experiments jointly support BR-MIL’s budgeted view: approximation improves with increasing $K$ (Theorem 5.1, Fig. 2) and the dominant term of the expensive relational component is governed by $K$ rather than $n$ (Theorem 5.2, Fig. 5).
Impact Statement
This paper presents work whose goal is to advance machine learning for budget-constrained relational multi-instance prediction, motivated by functional miRNA–mRNA target prediction. Potential positive impacts include improving the prioritization of candidate regulatory interactions for downstream experimental validation and enabling more compute-efficient analysis of large candidate pools. We do not anticipate major direct negative societal impacts of this work beyond standard risks of dataset bias and over-interpretation of model outputs.
Appendix A Extended Problem Setup and BR-MIL Formalization
A.1 Biological instantiation: miRNA–mRNA functional targeting
Task.
We study functional miRNA–mRNA targeting. Each example corresponds to a miRNA sequence $\mu^{(m)}$ and an mRNA $3^{\prime}\mathrm{UTR}$$\nu^{(m)}$, with a transcript-level label $Y^{(m)}\in\{0,1\}$ indicating whether $\mu^{(m)}$ functionally represses $\nu^{(m)}$. (We use $\mu,\nu$ to avoid overloading $u_{i}$ in the abstract formulation.)
Candidate target sites (CTSs).
Given $(\mu^{(m)},\nu^{(m)})$, we extract a variable-size set of candidate target sites (CTSs) from the 3’UTR, e.g., by sliding windows plus seed/alignment-based filtering:
$\mathcal{C}^{(m)}=\{c^{(m)}_{1},\dots,c^{(m)}_{n^{(m)}}\},\qquad n^{(m)}:=|\mathcal{C}^{(m)}|.$
Each CTS corresponds to a local window on $\nu^{(m)}$ that is potentially bindable by $\mu^{(m)}$. The candidate pool size $n^{(m)}$ varies substantially across pairs.
MIL view and mapping to our abstract bags.
We cast transcript-level prediction as multi-instance learning (MIL): the $m$-th miRNA–mRNA pair forms a bag of CTS instances. Concretely, each CTS $c^{(m)}_{i}$ induces (i) an instance content $x_{i}\in\mathcal{X}$ (e.g., the CTS window sequence/context features defined by $\mu^{(m)}$ and $\nu^{(m)}$), and (ii) structural attributes $u_{i}\in\mathcal{U}$ (e.g., normalized transcript position/region, seed type, alignment quality). This yields the abstract bag
$B^{(m)}=\{(x_{i},u_{i})\}_{i=1}^{n^{(m)}},$
with bag label $Y^{(m)}$ and latent instance labels (whether a CTS is truly functional). We next formalize this setting as Budgeted Relational MIL (BR-MIL) under a strict per-bag budget $K$.
A.2 Notation
We observe a dataset of bags
$\mathcal{D}=\{(B^{(m)},Y^{(m)},\mathcal{L}^{(m)})\}_{m=1}^{M}.$
Each bag $B^{(m)}$ corresponds to one miRNA–mRNA pair $(\mu^{(m)},\nu^{(m)})$ introduced in Sec. A.1, and its instances are the candidate target sites (CTSs) extracted from $\nu^{(m)}$ for $\mu^{(m)}$.
Each bag is a (multi)set of instances
$B=\{(x_{i},u_{i})\}_{i=1}^{n},$
where $x_{i}\in\mathcal{X}$ denotes instance content (e.g., CTS windows) and $u_{i}\in\mathcal{U}$ denotes structural attributes (e.g., transcript position/region). The bag label is $Y\in\{0,1\}$ (or $Y\in\mathbb{R}$ for regression). Optionally, a subset of instances is labeled: $\mathcal{L}\subseteq[n]$, with instance labels $\{\tilde{y}_{i}\}_{i\in\mathcal{L}}$.
[…]
Pairwise relations among selected instances may be provided explicitly as $R_{S}=\{r_{ij}\}_{i,j\in S}$ with
$r_{ij}=\rho(h_{i},h_{j},u_{i},u_{j}),$
or modeled implicitly by the aggregator. A permutation-invariant relational aggregator $f_{\phi}$ outputs the bag prediction
$\hat{Y}=f_{\phi}\!\big(\{(h_{i},u_{i})\}_{i\in S},\,R_{S}\big).$ (6)
When instance labels are available, an instance head $c_{\omega}$ yields $\hat{y}_{i}=c_{\omega}(h_{i})$ for $i\in S\cap\mathcal{L}$.
Definition A.1 (Budgeted Relational MIL (BR-MIL)).
Budgeted Relational MIL (BR-MIL) is a supervised learning problem characterized by: (i) large candidate pools ($n$ per bag), (ii) a strict per-bag budget $K$ on expensive encoding and relational aggregation, and (iii) modeling interaction effects among selected instances.
Formally, the hypothesis class consists of triplets $(\tilde{e}_{\tilde{\theta}},\pi_{\psi},f_{\phi})$ with prediction
$\hat{Y}(B)$ $=f_{\phi}\!\Big(\{(e_{\theta}(x_{i},u_{i}),u_{i})\}_{i\in S},\,R_{S}\Big),$ (7)
$S$ $\sim\pi_{\psi}\!\Big(\cdot\ \Big|\ \{\tilde{e}_{\tilde{\theta}}(x_{i},u_{i})\}_{i=1}^{n}\Big),\qquad|S|\leq K.$
Training minimizes the expected risk with mixed supervision:
$\min_{\tilde{\theta},\psi,\phi,\omega}$ (8)
$\mathbb{E}_{(B,Y,\mathcal{L})\sim\mathcal{D}}\ \mathbb{E}_{S\sim\pi_{\psi}(\cdot\mid\tilde{H})}\Big[\ell_{\text{bag}}(Y,\hat{Y})+\lambda\!\!\sum_{i\in S\cap\mathcal{L}}\!\!\ell_{\text{inst}}(\tilde{y}_{i},\hat{y}_{i})\Big]$
$\quad+\Omega(\pi_{\psi}),$
where $\Omega$ regularizes selection (e.g., diversity or exploration). The budget constraint is hard: expensive encoding and relational aggregation are executed only on $S$.
A.3 Permutation Invariance
A BR-MIL predictor is bag-permutation invariant if for any permutation $\pi$ of indices,
$\hat{Y}\big(\{(x_{\pi(i)},u_{\pi(i)})\}_{i=1}^{n}\big)=\hat{Y}\big(\{(x_{i},u_{i})\}_{i=1}^{n}\big).$ (9)
Sufficient conditions are: (i) the selector depends only on the multiset $\tilde{H}$ (not on order) and outputs $S$ via scores with either no ties or a fixed, data-independent tie-breaking rule; (ii) the aggregator $f_{\phi}$ is permutation invariant w.r.t. its selected inputs (e.g., DeepSets or Set Transformer); (iii) if explicit relations $R_{S}$ are constructed, the relation function $\rho$ is permutation-equivariant (i.e., permuting selected indices permutes $R_{S}$ consistently). In particular, Top-$K$ selection preserves invariance under (i)–(iii) because the prediction depends only on the set $S$.
A.4 Computational Complexity
For each bag, cheap encoding costs $\mathcal{O}(n)$ and is cacheable; selection costs $\mathcal{O}(n)$ (up to additional diversity heuristics). Expensive encoding costs $\mathcal{O}(K)$ and relational aggregation typically costs $\mathcal{O}(K^{2})$. Thus the deployed cost per bag is
$\mathcal{O}(n)+\mathcal{O}(K^{2}),$
under a strict access budget $K$, versus $\mathcal{O}(n^{2})$ if all instances were processed relationally.
Our proposed PAIR-Former instantiates this BR-MIL hypothesis class with (i) a TargetNet-compatible expensive encoder, (ii) a distilled cheap encoder for full-pool scanning, (iii) a deterministic budgeted selector (STSelector), and (iv) a Set Transformer aggregator.
Appendix B Full Theoretical Assumptions and Proofs
B.1 Preliminaries: masking view of budgeted execution
We formalize the “masking” interpretation used throughout the paper. Fix a bag $B=\{(x_{i},u_{i})\}_{i=1}^{n}$ and define the expensive token of instance $i$ as $z_{i}:=z(x_{i},u_{i})\in\mathbb{R}^{d_{z}},$ where $z(\cdot)$ denotes the composition of the expensive encoder and any tokenization (e.g., Eq. (4)). Let $Z(B):=(z_{1},\dots,z_{n})$ denote the (conceptual) full token list.
Reference full-information predictor.
Let $f_{\phi}$ be a permutation-invariant relational aggregator operating on a multiset of tokens. The reference predictor (conceptually) accesses all $n$ expensive tokens:
$\hat{Y}_{\mathrm{full}}(B):=f_{\phi}\big(\{z_{i}\}_{i=1}^{n}\big).$ (10)
Budgeted predictor via masking.
A (possibly data-dependent) selector outputs a subset $S(B)\subseteq[n]$ with $|S(B)|\leq K$. Fix a padding token $z_{\emptyset}\in\mathbb{R}^{d_{z}}$ and define the masked token list $Z^{\mathrm{mask}(S)}(B):=(z_{1}^{\prime},\dots,z_{n}^{\prime})$ by $z_{i}^{\prime}=z_{i}$ if $i\in S$ and $z_{i}^{\prime}=z_{\emptyset}$ if $i\notin S$. The budgeted prediction is
$\hat{Y}_{K}(B):=f_{\phi}\big(\{z_{i}^{\prime}\}_{i=1}^{n}\big)=f_{\phi}\big(\{z_{i}\}_{i\in S(B)};\;\text{mask}\big).$ (11)
In practice (Set Transformer / SAB), the “mask” is implemented by attention masks so that padded tokens do not contribute to attention weights or pooling, making (11) a faithful abstraction.
B.2 Formal approximation bound (Theorem 5.1)
We state a fully formal version of Theorem 5.1. The key object is an influence weight distribution $w(B)=(w_{1}(B),\dots,w_{n}(B))$ over instances, which quantifies how sensitive the aggregator output is to each token under masking.
Assumption B.1 (Mask-consistent execution).
There exists a fixed padding token $z_{\emptyset}$ and a masking convention such that the implemented budgeted forward pass (padding to kmax and masking in attention/pooling) is equivalent to evaluating $f_{\phi}$ on the masked token multiset $\{z_{i}^{\prime}\}_{i=1}^{n}$ defined above.
Assumption B.2 (Bounded token radius).
There exists $R>0$ such that for all bags and all instances, $\|z_{i}\|_{2}\leq R$ and $\|z_{\emptyset}\|_{2}\leq R$.
Assumption B.3 (Weighted Lipschitz stability under masking).
For each bag $B$, there exist weights $w_{i}(B)\geq 0$ with $\sum_{i=1}^{n}w_{i}(B)=1$ and a constant $L_{f}>0$ such that for any subset $S\subseteq[n]$,
$|\hat{Y}_{\mathrm{full}}(B)-f_{\phi}(Z^{\mathrm{mask}(S)}(B))|\;\leq\;L_{f}\sum_{i\notin S}w_{i}(B)\,\|z_{i}-z_{\emptyset}\|_{2}.$ (12)
A sufficient condition for Assumption B.3.
Assumption B.3 is mild: it holds whenever the aggregator is differentiable and the per-token gradient norms are controlled.
Lemma B.4 (Gradient-based influence weights imply (12)).
Assume $f_{\phi}$ is differentiable in each token. Let $\mathcal{Z}(B)$ denote the set of all masked states obtained by replacing any subset of tokens by $z_{\emptyset}$. Define per-token sensitivities
$g_{i}(B):$ $=\sup_{Z\in\mathcal{Z}(B)}\big\|\nabla_{z_{i}}f_{\phi}(Z)\big\|_{2},$
$G(B):$ $=\sum_{j=1}^{n}g_{j}(B),$
$w_{i}(B):$ $=\frac{g_{i}(B)}{G(B)}.$
Then for any subset $S\subseteq[n]$,
$\Big|f_{\phi}(Z(B))-f_{\phi}\big(Z^{\mathrm{mask}(S)}(B)\big)\Big|\leq\sum_{i\notin S}g_{i}(B)\,\|z_{i}-z_{\emptyset}\|_{2}$
$=G(B)\sum_{i\notin S}w_{i}(B)\,\|z_{i}-z_{\emptyset}\|_{2},$
so (12) holds with $L_{f}:=G(B)$.
Proof.
Let $Z^{(0)}:=Z(B)$ and obtain $Z^{(t)}$ by masking one additional token at each step until reaching $Z^{\mathrm{mask}(S)}(B)$ (mask exactly those $i\notin S$), so that only one token changes per step. By the mean value theorem applied to the $i$-th token at step $t$,
$\big|f_{\phi}(Z^{(t-1)})-f_{\phi}(Z^{(t)})\big|$ $\leq\sup_{Z\in\mathcal{Z}(B)}\|\nabla_{z_{i}}f_{\phi}(Z)\|_{2}\,\|z_{i}-z_{\emptyset}\|_{2}$
$=g_{i}(B)\,\|z_{i}-z_{\emptyset}\|_{2}.$
Summing over all masked tokens $i\notin S$ yields the bound. ∎
Assumption B.5 (Selector covers influence mass).
For budget $K$, the selector output $S(B)$ satisfies
$\mathbb{E}\Big[\sum_{i\in S(B)}w_{i}(B)\ \Big|\ B\Big]\;\geq\;1-\varepsilon_{K}(B),$ (13)
where the expectation is over any randomness in the selector (and equals identity for deterministic selection).
Assumption B.6 (Tail mass + selector suboptimality (optional rate form)).
Let $w_{(1)}(B)\geq w_{(2)}(B)\geq\cdots$ be the weights sorted in descending order and define the tail
$\psi(K;B):=\sum_{i>K}w_{(i)}(B).$ (14)
Define the selector suboptimality at budget $K$ as
$\delta_{K}(B):=\Big(\sum_{i=1}^{K}w_{(i)}(B)\Big)-\mathbb{E}\Big[\sum_{i\in S(B)}w_{i}(B)\ \Big|\ B\Big]\ \geq 0.$ (15)
Then $\varepsilon_{K}(B)\leq\psi(K;B)+\delta_{K}(B)$. Optionally, for an explicit decay rate one may assume a heavy-tail bound such as $\psi(K;B)\leq CK^{1-\alpha}$ for some $\alpha>1$ and bag-dependent constant $C$.
Lemma B.7 (Best-$K$ mass equals the sorted top-$K$ sum).
For any nonnegative weights $w_{i}$ with $\sum_{i}w_{i}=1$,
$\max_{|S|\leq K}\ \sum_{i\in S}w_{i}\;=\;\sum_{i=1}^{K}w_{(i)}.$
Consequently, the minimal uncovered mass achievable by any size-$K$ subset equals $\psi(K)$.
Proof.
The maximum is achieved by selecting the $K$ largest weights; any other subset can be improved by exchanging a smaller selected weight with a larger unselected weight. The uncovered mass is $1-\sum_{i\in S}w_{i}$, so its minimum is $1-\sum_{i=1}^{K}w_{(i)}=\psi(K)$. ∎
Theorem B.8 (Formal approximation bound under budgeted masking).
Assume Assumptions B.1B.5. Then for any fixed bag $B$,
$\mathbb{E}\Big[\,\big|\hat{Y}_{\mathrm{full}}(B)-\hat{Y}_{K}(B)\big|\,\Big|\ B\Big]\;\leq\;2L_{f}R\cdot\varepsilon_{K}(B),$ (16)
where the expectation is over selector randomness. If Assumption B.6 also holds, then
$\mathbb{E}\Big[\,\big|\hat{Y}_{\mathrm{full}}(B)-\hat{Y}_{K}(B)\big|\,\Big|\ B\Big]\;\leq\;2L_{f}R\cdot\big(\psi(K;B)+\delta_{K}(B)\big),$ (17)
and under $\psi(K;B)\leq CK^{1-\alpha}$ this yields a polynomial decay in $K$ (up to $\delta_{K}$). Moreover, taking expectation over $B\sim\mathcal{P}$ gives the population version: $\mathbb{E}_{B}\mathbb{E}[\ |\hat{Y}_{\mathrm{full}}(B)-\hat{Y}_{K}(B)|\mid B]\leq 2L_{f}R\,\mathbb{E}_{B}[\varepsilon_{K}(B)]$.
Proof.
By Assumption B.1 and the definition of $\hat{Y}_{K}$ in (11),
$\hat{Y}_{K}(B)=f_{\phi}(\{z_{i}^{\prime}\}_{i=1}^{n})\quad\text{with}\quad z_{i}^{\prime}=\begin{cases}z_{i},&i\in S(B),\\ z_{\emptyset},&i\notin S(B).\end{cases}$
Applying Assumption B.3 to the subset $S(B)$ gives
$\big|\hat{Y}_{\mathrm{full}}(B)-\hat{Y}_{K}(B)\big|\leq L_{f}\sum_{i\notin S(B)}w_{i}(B)\,\|z_{i}-z_{\emptyset}\|_{2}.$
By Assumption B.2, $\|z_{i}-z_{\emptyset}\|_{2}\leq\|z_{i}\|_{2}+\|z_{\emptyset}\|_{2}\leq 2R$, hence
$\big|\hat{Y}_{\mathrm{full}}(B)-\hat{Y}_{K}(B)\big|\leq 2L_{f}R\sum_{i\notin S(B)}w_{i}(B)$
$=2L_{f}R\Big(1-\sum_{i\in S(B)}w_{i}(B)\Big).$
Taking conditional expectation over selector randomness and using Assumption B.5 yields (16). Under Assumption B.6, we have $\varepsilon_{K}(B)\leq\psi(K;B)+\delta_{K}(B)$, giving (17). The population statement follows by taking expectation over $B$. ∎
Connection to Theorem 5.1 in the main text.
Theorem 5.1 is the informal/summary form of Theorem B.8: it emphasizes that the truncation/masking gap is controlled by the uncovered influence mass (and thus decreases as $K$ increases and/or the selector reduces $\delta_{K}$).
B.3 Generalization controlled by the budget $K$ (Theorem 5.2)
We provide a formal generalization result for the expensive relational component in BR-MIL. The key message is that once expensive computation is restricted to at most $K$ selected tokens per bag, the dominant capacity term of the expensive relational predictor depends on $K$ (and token dimension), not on the raw pool size $n$. Any additional $n$-dependence can only enter through the selector family.
Setup.
Let $\mathcal{P}$ be a distribution over labeled bags $(B,Y)$. We observe $M$ i.i.d. samples $\mathcal{D}=\{(B^{(m)},Y^{(m)})\}_{m=1}^{M}$. For each bag $B$, a selector $S(\cdot)$ outputs $S(B)\subseteq[n]$ with $|S(B)|\leq K$. Let the corresponding selected expensive tokens be
$Z_{S(B)}:=\{z_{i}(B)\}_{i\in S(B)}.$
For analysis, we view $Z_{S(B)}$ as a padded matrix in $\mathbb{R}^{K\times d_{z}}$ (by appending $z_{\emptyset}$ tokens as needed) and denote its vectorization by
$\mathrm{vec}(Z_{S(B)})\in\mathbb{R}^{Kd_{z}}.$
Assumption B.9 (i.i.d. bags).
$(B^{(m)},Y^{(m)})$ are drawn i.i.d. from $\mathcal{P}$.
Assumption B.10 (Bounded selected tokens).
There exists $R>0$ such that for any bag $B$ and any $i\in S(B)$, $\|z_{i}(B)\|_{2}\leq R$ (and $\|z_{\emptyset}\|_{2}\leq R$). Equivalently, $\|\mathrm{vec}(Z_{S(B)})\|_{2}\leq R\sqrt{K}$ for all bags.
Assumption B.11 (Lipschitz loss).
The loss $\ell:\mathcal{Y}\times\mathbb{R}\to[0,1]$ is $L_{\ell}$-Lipschitz in its prediction argument: for all $y$ and all $a,b\in\mathbb{R}$, $|\ell(y,a)-\ell(y,b)|\leq L_{\ell}|a-b|$.
Assumption B.12 (Lipschitz aggregator class on $K$ tokens).
Let $\mathcal{F}_{K}$ be a class of (permutation-invariant) aggregators that operate on at most $K$ tokens (e.g., a Set Transformer of fixed depth/width with masking). Assume each $f\in\mathcal{F}_{K}$ is $L_{f}$-Lipschitz w.r.t. the Frobenius norm on padded token matrices: for all $A,B\in\mathbb{R}^{K\times d_{z}}$,
$\qquad\qquad\qquad|f(A)-f(B)|\leq L_{f}\|A-B\|_{F}$
$\begin{aligned} (\text{equivalently }|f(\mathrm{vec}(A))&-f(\mathrm{vec}(B))|\leq L_{f}\|\mathrm{vec}(A)\\ &-\mathrm{vec}(B)\|_{2}).\end{aligned}$
Assumption B.13 (Selector family).
Selectors belong to a family $\mathcal{S}_{K}$ such that for all $S\in\mathcal{S}_{K}$ and all bags $B$, $|S(B)|\leq K$. When we need a uniform bound over selectors, we assume $\mathcal{S}_{K}$ is finite with $|\mathcal{S}_{K}|=N_{S}$.
Risk.
For $(f,S)\in\mathcal{F}_{K}\times\mathcal{S}_{K}$, define the predictor $\hat{y}(B)=f(Z_{S(B)})$ and the loss function
$h_{f,S}(B,Y):=\ell\big(Y,f(Z_{S(B)})\big).$
Let the population and empirical risks be
$\mathcal{R}(f,S):$ $=\mathbb{E}_{(B,Y)\sim\mathcal{P}}\big[h_{f,S}(B,Y)\big],$
$\widehat{\mathcal{R}}(f,S):$ $=\frac{1}{M}\sum_{m=1}^{M}h_{f,S}(B^{(m)},Y^{(m)}).$
Rademacher complexity.
For a class $\mathcal{H}$ of functions mapping examples to $[0,1]$, define its empirical Rademacher complexity as
$\widehat{\mathfrak{R}}_{M}(\mathcal{H}):=\mathbb{E}_{\varepsilon}\Big[\sup_{h\in\mathcal{H}}\frac{1}{M}\sum_{m=1}^{M}\varepsilon_{m}\,h(B^{(m)},Y^{(m)})\Big],$
where $\varepsilon_{1},\dots,\varepsilon_{M}$ are i.i.d. Rademacher variables. Let $\mathfrak{R}_{M}(\mathcal{H}):=\mathbb{E}_{\mathcal{D}}[\widehat{\mathfrak{R}}_{M}(\mathcal{H})]$.
We consider two classes:
$\mathcal{H}_{K,S}:=\{(B,Y)\mapsto\ell(Y,f(Z_{S(B)})):\ f\in\mathcal{F}_{K}\}~(\text{$S$ fixed}),$
and the union over selectors
$\mathcal{H}_{K}=\bigcup_{S\in\mathcal{S}_{K}}\mathcal{H}_{K,S}$ $=\{(B,Y)\mapsto\ell(Y,f(Z_{S(B)})):$
$f\in\mathcal{F}_{K},\,S\in\mathcal{S}_{K}\}.$
Theorem B.14 (Generalization governed by budget $K$; formal version).
Assume B.9, B.10, B.11, B.12 and B.17. Fix any selector $S$ (deterministic or conditioned upon), and consider $\mathcal{H}_{K,S}$. Then for any $\delta\in(0,1)$, with probability at least $1-\delta$ over the draw of $\mathcal{D}$, for all $f\in\mathcal{F}_{K}$,
$\mathcal{R}(f,S)\;\leq\;\widehat{\mathcal{R}}(f,S)+2\,\mathfrak{R}_{M}(\mathcal{H}_{K,S})+3\sqrt{\frac{\log(2/\delta)}{2M}}.$ (18)
Moreover, the Rademacher complexity of the expensive relational component satisfies
$\mathfrak{R}_{M}(\mathcal{H}_{K,S})\;\leq\;C\,\frac{L_{\ell}L_{f}R}{\sqrt{M}}\,\sqrt{K},$ (19)
for a universal constant $C$ (absorbing fixed token dimension factors). Thus the dominant term scales as $\tilde{\mathcal{O}}(\sqrt{K/M})$ and is independent of the raw pool size $n$.
If additionally B.13 holds with $|\mathcal{S}_{K}|=N_{S}<\infty$, then with probability at least $1-\delta$, uniformly for all $(f,S)\in\mathcal{F}_{K}\times\mathcal{S}_{K}$,
$\mathcal{R}(f,S)\;\leq\;\widehat{\mathcal{R}}(f,S)+C\,\frac{L_{\ell}L_{f}R}{\sqrt{M}}\,\sqrt{K}+3\sqrt{\frac{\log(2N_{S}/\delta)}{2M}}.$ (20)
Interpretation (where $n$ can enter).
The bound isolates that the expensive relational component depends on $K$ (and fixed token dimension), not on $n$. Any $n$-dependence can only enter through the selector family size $N_{S}$. For instance, if $\mathcal{S}_{K}$ contains all $K$-subsets of an $n$-candidate pool, then $\log N_{S}=\log\sum_{j=0}^{K}{n\choose j}=\mathcal{O}(K\log(en/K))$, whereas for a deterministic Top-$K$ rule with fixed tie-breaking, $N_{S}=1$ and this term vanishes.
Lemma B.15 (Standard Rademacher generalization bound).
For any $\mathcal{H}$ of functions into $[0,1]$ and any $\delta\in(0,1)$, with probability at least $1-\delta$ over $\mathcal{D}$,
$\forall h\in\mathcal{H}:\quad\mathbb{E}_{(B,Y)\sim\mathcal{P}}\big[h(B,Y)\big]\leq\frac{1}{M}\sum_{m=1}^{M}h\big(B^{(m)},Y^{(m)}\big)$
$+2\,\mathfrak{R}_{M}(\mathcal{H})+3\sqrt{\frac{\log(2/\delta)}{2M}}.$
Proof.
This is a standard consequence of symmetrization, concentration (Hoeffding), and the definition of Rademacher complexity (e.g., Bartlett–Mendelson style bounds). ∎
Lemma B.16 (Contraction for Lipschitz losses).
Assume B.11. Let $\mathcal{G}$ be a class of real-valued functions on bags, and define $\ell\circ\mathcal{G}:=\{(B,Y)\mapsto\ell(Y,g(B)):\ g\in\mathcal{G}\}$. Then
$\mathfrak{R}_{M}(\ell\circ\mathcal{G})\leq L_{\ell}\,\mathfrak{R}_{M}(\mathcal{G}).$
Proof.
This is the standard contraction inequality for Rademacher averages applied pointwise in $Y$, using the $L_{\ell}$-Lipschitz property in the prediction argument. ∎
Assumption B.17 (Capacity control of $\mathcal{F}_{K}$).
Fix any selector $S$ and let
$\mathcal{G}_{K,S}:=\{B\mapsto f(Z_{S(B)}):\ f\in\mathcal{F}_{K}\}.$
Assume there exists a constant $C_{0}$ (depending on norm/architecture control of $\mathcal{F}_{K}$ but not on $n$) such that
$\mathfrak{R}_{M}(\mathcal{G}_{K,S})\leq C_{0}\,\frac{L_{f}R}{\sqrt{M}}\,\sqrt{K}.$
Proof of Theorem B.14.
Apply Lemma B.15 to $\mathcal{H}_{K,S}$ to obtain (18). For the complexity term, write $\mathcal{H}_{K,S}=\ell\circ\mathcal{G}_{K,S}$. By Lemma B.16, $\mathfrak{R}_{M}(\mathcal{H}_{K,S})\leq L_{\ell}\mathfrak{R}_{M}(\mathcal{G}_{K,S})$, and by B.17, $\mathfrak{R}_{M}(\mathcal{G}_{K,S})\leq C_{0}(L_{f}R/\sqrt{M})\sqrt{K}$. Combining yields (19) with $C=C_{0}$ (or another universal constant).
For the uniform bound over $(f,S)$ when $|\mathcal{S}_{K}|=N_{S}<\infty$, apply Lemma B.15 to each class $\mathcal{H}_{K,S}$ with failure probability $\delta/N_{S}$. By a union bound over $S\in\mathcal{S}_{K}$, with probability at least $1-\delta$, simultaneously for all $S\in\mathcal{S}_{K}$ and all $f\in\mathcal{F}_{K}$,
$\mathcal{R}(f,S)\leq\widehat{\mathcal{R}}(f,S)+2\,\mathfrak{R}_{M}(\mathcal{H}_{K,S})+3\sqrt{\frac{\log(2N_{S}/\delta)}{2M}}.$
Using the fixed-selector complexity bound (19) and absorbing constants into $C$ yields (20).
Appendix C Algorithms and Pseudocode
Algorithm 1 Three-stage training for BR-MIL on miRNA–mRNA targeting
1:Stage 1 (expensive CTS encoder). Train $e_{\theta}$ on miRNA–CTS pairs with binary loss (Sec. 4.7).
2:Stage 2 (cheap CTS encoder). Train $\tilde{e}_{\tilde{\theta}}$ by distilling from $e_{\theta}$ using Eq. (5).
3:Stage 3 (aggregator + joint fine-tune). For each miRNA–mRNA pair $(\mu,\nu)$:
4:  Extract CTS candidates via ESA scanning/filtering $\Rightarrow\{(x_{i},u_{i})\}_{i=1}^{n}$.
5:  Cheap scan: compute $(\tilde{h}_{i},\tilde{z}_{i})=\tilde{e}_{\tilde{\theta}}(x_{i},u_{i})$ for all $i\in[n]$.
6:  Select $S$ via STSelector on $\{(\tilde{h}_{i},\tilde{z}_{i},p_{i})\}_{i=1}^{n}$ with budget $K=\min(\texttt{kmax},n)$.
7:  Expensive encode only selected: $(h_{i},z_{i})=e_{\theta}(x_{i},u_{i})$ for $i\in S$.
8:  Tokenize: $t_{i}=[h_{i}\|z_{i}\|s_{i}^{\mathrm{esa}}\|p_{i}]$; pad/mask to kmax.
9:  Aggregate: $z_{\text{pair}}=f_{\phi}(\{t_{i}\}_{i\in S})$; compute $L_{\text{pair}}=\mathrm{BinaryLoss}(z_{\text{pair}},Y)$.
10:  Warmup: freeze $\theta$, update only $\phi$; Joint FT: unfreeze $\theta$, update $(\theta,\phi)$ for a few epochs.
Algorithm 2 Budgeted inference for one miRNA–mRNA pair
1:Input: $(\mu,\nu)$, cheap encoder $\tilde{e}_{\tilde{\theta}}$, selector, expensive encoder $e_{\theta}$, aggregator $f_{\phi}$, kmax=64
2: Extract CTS candidates by ESA scan/filter: $\{(x_{i},u_{i})\}_{i=1}^{n}$ with $s_{i}^{\mathrm{esa}}\geq 6$.
3: Compute $(\tilde{h}_{i},\tilde{z}_{i})=\tilde{e}_{\tilde{\theta}}(x_{i},u_{i})$ for all $i\in[n]$.
4: Select $S$ with $|S|=K=\min(\texttt{kmax},n)$ via STSelector (Top-$K_{1}$ + diversity).
5: Compute $(h_{i},z_{i})=e_{\theta}(x_{i},u_{i})$ for $i\in S$.
6: Form tokens $t_{i}=[h_{i}\|z_{i}\|s_{i}^{\mathrm{esa}}\|p_{i}]$, pad/mask to kmax.
7: Output $z_{\text{pair}}=f_{\phi}(\{t_{i}\}_{i\in S})$ and $\hat{Y}=\sigma(z_{\text{pair}})$.
Appendix D Additional Details for STSelector
This appendix provides full algorithmic details of STSelector, which is summarized in the main text as Top-$K$ exploitation + position coverage + embedding deduplication.
Inputs.
Given a bag with cheap embeddings $\{\tilde{h}_{i}\}_{i=1}^{n}$, cheap logits $\{\tilde{z}_{i}\}_{i=1}^{n}$, and normalized transcript positions $\{p_{i}\}_{i=1}^{n}$, the selector outputs a subset $S\subseteq[n]$ with $|S|\leq K$.
Step A (Top-$K_{1}$ exploitation).
We select
$S_{1}=\mathrm{TopK}(\tilde{z}_{i},K_{1}),$
favoring highly confident candidates under the cheap encoder.
Step B (Position binning).
We partition candidates into $B$ bins according to transcript position $p_{i}\in[0,1]$. Within each bin, we keep a heap of size $m$ containing the top-$m$ candidates by $\tilde{z}_{i}$. This produces a reduced pool $C$ of size at most $B\cdot m$.
Step C (Embedding deduplication).
To reduce redundancy, we compute a lightweight SimHash key on $\tilde{h}_{i}$ (sign bits of selected dimensions). Within each bin, at most $c$ candidates are kept per hash key.
Step D (Balanced quota allocation).
For each bin $b$, we compute a weight
$w_{b}=\sum_{i\in\mathrm{Top-}t(b)}\exp(\tilde{z}_{i}/\tau_{w}),$
where $\mathrm{Top-}t(b)$ denotes the top-$t$ candidates in bin $b$. We allocate a quota of candidates to each bin proportional to $w_{b}$, enforcing at least one per bin when possible.
Step E (Merge and fill).
We output
$S=\mathrm{dedup}(S_{1}\cup S_{2}),$
and fill remaining slots by descending $\tilde{z}_{i}$ until $|S|=K$.
Complexity.
STSelector runs in $\mathcal{O}(n\log m)$ time due to per-bin heaps, and is implemented entirely on CPU for low-latency inference.
Appendix E Loss Functions and Distillation Details
Binary loss.
Given logit $z$ and label $y\in\{0,1\}$, we apply label smoothing
$\tilde{y}=\begin{cases}0.95&y=1,\\ 0.05&y=0.\end{cases}$
We use weighted BCE loss
$\ell_{\mathrm{BCE}}(z,\tilde{y})=-w\big[\tilde{y}\log\sigma(z)+(1-\tilde{y})\log(1-\sigma(z))\big].$
Focal mixture.
We optionally combine BCE with focal reweighting:
$\ell_{\mathrm{focal}}=\alpha_{t}(1-p_{t})^{\gamma}\ell_{\mathrm{BCE}},$
with $\gamma=1$, $\alpha=0.4$. Final loss is
$L=\lambda_{\mathrm{bce}}L_{\mathrm{BCE}}+\lambda_{\mathrm{focal}}L_{\mathrm{focal}},\quad(\lambda_{\mathrm{bce}},\lambda_{\mathrm{focal}})=(0.01,1).$
Distillation.
The cheap encoder is trained via
$\mathcal{L}_{\mathrm{distill}}=(1-\alpha)\mathcal{L}_{\mathrm{sup}}+\alpha\,\mathcal{L}_{\mathrm{KD}}+\beta_{\mathrm{feat}}\,\mathcal{L}_{\mathrm{feat}}+\beta_{\mathrm{rel}}\,\mathcal{L}_{\mathrm{rel}}.$ (21)
with temperature $T=2$ and cosine schedule $\alpha:0.8\to 0.5$. We use $(\beta_{\mathrm{feat}},\beta_{\mathrm{rel}})=(0.1,1)$.
Appendix F Full Metrics for Main Comparison
Table 3 reports additional derived metrics omitted from the main table for clarity.
Table 3: Additional metrics for main comparison (miRAW half-split).
Method Acc Prec Rec Spec
TargetNet (reproduced) 0.7226 0.6560 0.9361 0.5091
Max pooling (ours) 0.8049 0.7822 0.8453 0.7646
PAIR-Former (ours) 0.9744 0.9886 0.9598 0.9890
Appendix G Additional Related Work and Biological Context
From strongest-site pooling to bag-level reasoning.
Many functional predictors adopt an OR-like transcript rule implemented as max pooling over site scores (Min et al., 2022; Pla and others, 2018; Li and others, 2024; Yang et al., 2024). While effective and simple, max pooling hard-codes a single-dominant-instance assumption that can be misaligned with transcripts containing multiple plausible CTSs, especially under relaxed candidate generation where redundancy is common (Min et al., 2022; Sætrom et al., 2007).
Evidence for cooperative/competitive effects.
Prior biological studies emphasize that repression can depend on the number of sites, their spacing, and local context (Bartel, 2009; Sætrom et al., 2007). Moreover, competition via alternative targets or decoy-like binding sites can modulate apparent transcript-level outcomes, weakening the assumption that one strongest site fully explains repression (Salmena et al., 2011). These observations motivate MIL formulations where the prediction depends on a set of latent instance events rather than a single site.
Appendix H Additional runtime profiling details
Goal and setting.
This appendix documents the profiling protocol behind Figures 3 and 4. We measure online inference cost for three pipelines: BR-MIL_online (cheap scan + STSelector + expensive encode on $K$ + Set Transformer), TargetNet_like_online (window/CTS scoring + pooling; budget-independent), and a heavier Naive_online variant that performs more per-candidate computation before aggregation.
Hardware and software.
All measurements are conducted on a single NVIDIA RTX 4090 (24 GB). We use PyTorch 2.4.1 (+cu121) with CUDA 12.1 and cuDNN 9.1.0. CPU model, number of threads, and OS details are reported in our released supplementary code/configs.
Measurement protocol.
[…]
Definition of stages in Figure 4.
The stage breakdown aggregates profiled blocks that appear in all pipelines:
  • CTS generation / filtering: ESA scan over the 3UTR and candidate filtering.
  • CPU gather (gather_all_cpu): CPU-side packing/gathering of candidate tensors and metadata into contiguous buffers for subsequent model calls.
  • Scan / cheap forward (scan_or_cheap_forward): per-candidate evaluation of the cheap encoder (or the corresponding per-site scorer in non-BR-MIL baselines).
  • Selection (STSelector): CPU-only subset selection producing indices $S$ with $|S|=K$.
  • Expensive encode on $K$: forward pass of the expensive CTS encoder applied only to selected candidates.
  • Tokenize + pack: token concatenation (Eq. (4)), padding/masking to kmax, and device transfers if needed.
  • Aggregation: Set Transformer (SAB stack + PMA) producing the pair logit.
Stage names in the figures correspond to these blocks; minor implementation-level sub-stages are merged for readability.
Peak VRAM reporting.
Peak VRAM is measured as the maximum allocated GPU memory during the forward pass under the same input batch and $K$. Since TargetNet_like_online does not allocate token batches of size $K$ nor run attention over $K$ tokens, its VRAM usage is largely flat across the $K$ sweep, whereas budgeted relational pipelines scale with $K$ due to token packing and attention.
Absolute numbers and configuration table.
For completeness, we provide the full table of absolute latency/throughput/VRAM values (and any additional profiling knobs such as batch size, number of CPU workers/threads, and I/O caching settings) in the supplementary material code/configs released with this submission.